# Circumgalactic Enrichment by a Multi-Galaxy Merger at Redshift 6.7

### v1

@@ Methods
We estimate the merging timescales in three ways. We first estimate the lower limit of the merging timescale, assuming the secondary galaxy follows a circular orbit about the primary galaxy, where dynamical friction dominates its energy and angular momentum losses. The merging timescale follows the formula (Equation 12.45 in ref [84]) based on Chandrasekhar’s dynamical friction formula [85],
@@
Second, equation 1 can underestimate the merging timescale because, during the merger, the tidal forces of the primary galaxy would gradually strip the matter from the secondary galaxy, decreasing its mass [86]. This effect would increase the merging timescale by a factor of 2 – 5 [87, 84]. To incorporate the effect of mass loss, we adopt the formula by Ref [88] (Equation 8.17),
@@
Third, we adopt a the formula of ref [87] (Equation 8), which fits results from an -body numerical simulations, to estimate the merging timescale:
@@
Can outflows explain the [O III]+H bridge between ELG1 and ELG2?
Although the JWST NIRSpec spectrum of ELG1 does not reveal evidence of outflows, this does not conclusively rule out the presence of outflows in other clumps within ELG1 or in ELG2, which were not targeted by the NIRSpec observations. In this section, we further explore whether the outflows in ELG1 and ELG2 and their adjacent clumps can produce the [O III]+H bridge between them based on the outflow simulations. We estimate the [O III]+H luminosity of [O III]+H bridge to be the total flux enclosed by a polygon region defined in Extended Data Figure 7 with the contribution of clumps C12 and C13 removed based on Extended Data Table 1. We derive the total observed [O III]+H luminosity of the bridge to be erg s-1.
@@
Code availability Codes used in this study are publicly available: Astropy (https://www.astropy.org), Bagpipes (https://github.com/ACCarnall/bagpipes), CIGALE (https://cigale.lam.fr), Pysersic (https://github.com/pysersic/pysersic), sep (https://github.com/sep-developers/sep), and SExtractor (https://www.astromatic.net/software/sextractor).

### v2

@@ 1.9 Merging timescale of the JQ system
We estimate the merging timescales in three methods. In the first method, we estimate the lower limit of the merging timescale, assuming the secondary galaxy follows a circular orbit about the primary galaxy, where dynamical friction dominates its energy and angular momentum losses. The merging timescale follows the formula (Equation 12.45 in ref [94]) based on Chandrasekhar’s dynamical friction formula [95],
@@
However, equation 1 can underestimate the merging timescale because, during the merger, the tidal forces of the primary galaxy would gradually strip the matter from the secondary galaxy, decreasing its mass [96]. This effect would increase the merging timescale by a factor of 2 – 5 [97, 94]. To incorporate the effect of mass loss, we adopt the formula by Ref [98] (Equation 8.17) as the second method,
@@
In the third method, we adopt a the formula of ref [97] (Equation 8), which fits results from an -body numerical simulations, to estimate the merging timescale:
@@ 1.10 No evidence of outflows in ELG1
1.11 Can outflows explain the [O III]+H bridge between ELG1 and ELG2?
Although the JWST NIRSpec spectrum of ELG1 does not reveal evidence of outflows, this does not conclusively rule out the presence of outflows in other clumps within ELG1 or in ELG2, which were not targeted by the NIRSpec observations. In this section, we further explore whether the outflows in ELG1 and ELG2 and their adjacent clumps can produce the [O III]+H bridge between them based on the outflow simulations. We estimate the [O III]+H luminosity of [O III]+H bridge to be the total flux enclosed by a polygonal region defined in Extended Data Figure 8 with the contribution of clumps C12 and C13 removed based on Extended Data Table 1. We derive the total observed [O III]+H luminosity of the bridge to be erg s-1.
@@ 1.11 Can outflows explain the [O III]+H bridge between ELG1 and ELG2?
Code availability Codes used in this study are publicly available: Astropy (https://www.astropy.org), Bagpipes (https://github.com/ACCarnall/bagpipes), CIGALE (https://cigale.lam.fr), Pysersic (https://github.com/pysersic/pysersic), sep (https://github.com/sep-developers/sep), SExtractor (https://www.astromatic.net/software/sextractor, and MVT binning (https://github.com/pierrethx/MVT-binning).